if sin theta +sin ^2 theta =1 prove that cos^2 +cos^4=1
Answer by mananth(16946)

sin theta +sin ^2 theta =1 prove that cos^2 +cos^4=1

sin (theta) +sin ^2 (theta) =1

sin (theta) +1-cos ^2 (theta) =1

sin (theta) -cos ^2 (theta) =0

sin (theta) =cos ^2 (theta)
square both sides
sin^2 (theta) =cos ^4 (theta)
1-cos ^2 (theta)= cos ^4 (theta)
cos ^4 (theta)+cos ^2 (theta)=1